# A central limit theorem for a generalization of the Ewens measure to random tuples of commuting permutations

### v1

@@ 1.1. The main result
For $n\geq 0$, the symmetric group $\mathfrak{S}_{n}$ comes with a natural left action $(\sigma,i)\mapsto\sigma(i)$ on the set $[n]:=\{1,2,\ldots,n\}$. If $H$ is a subgroup of $\mathfrak{S}_{n}$, then the restriction of this action to $H$ gives us access to the enumerative quantity $\kappa_{n}(H):=|[n]/H|$, where we used ||$\cdot$|| ⋅ | to denote the cardinality of finite sets, and where $[n]/H$ is the set of orbits for the action of $H$ on $[n]$. We will be particularly interested in the situation where $H=\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle$ is the subgroup generated by a random ordered $\ell$-tuple $(\sigma_{1},\ldots,\sigma_{\ell})$ of pairwise commuting permutations, for some fixed $\ell\geq 1$. Our main result is a central limit theorem for the number of joint orbits $\kappa_{n}(\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle)$ or simply $\kappa_{\ell,n}(\sigma_{1},\ldots,\sigma_{\ell})$. More precisely, let us introduce the set of commuting tuples of permutations
@@
We view $\mathbb{P}_{\ell,n,x}$ as a natural generalization of the Ewens measure with parameter $x$ to tuples of commuting permutations. Indeed, when $\ell=1$, $\kappa_{1,n}(\sigma)$ is just the number of cycles of the single permutation $\sigma$, and the well known expansion of the Pochammer symbol in terms of Stirling numbers of the first kind gives the following specialization of the above PMF formula
@@
This is the PMF of the much studied Ewens distribution on the symmetric group, which was introduced in the applied context of mathematical biology [13], but is also an important notion from a pure mathematics point of view (see [29, 11, 33] for insightful reviews, and the book [4] for an in-depth introduction to the Ewens measure and many other related topics in probability theory and combinatorics). The uniform $x=1$ case was of course studied much earlier than the article by Ewens, but we also note that the particular weight $x=2$ per cycle (in a more difficult spatially dependent rather than mean field situation) features in the work of Tóth on quantum spin systems [35, §5].
@@
proved by Hansen as the time 1 projection of a functional CLT for a process related to the Ewens measure [20]. For $\ell=1$, the mean and variance are both asymptotically equivalent to $x\ln n$. The $x=1$ uniform sampling case is much older and due to Goncharov [16, 17] and is often given as an instructive example in graduate probability textbooks (see, e.g., [6, Example 27.3]). Using the Feller coupling idea [14, p. 815], one can express $\mathsf{K}_{1,n}$ as a sum of $n$ independent Bernoulli random variables of parameters $\frac{1}{n-j+1}$, $1\leq j\leq n$. Since these are not identically distributed, they provide a nice pedagogical example of application of the Lindeberg CLT. We also note the remarkable continuity of the formulas for the asymptotic mean and variance of the $\mathsf{K}_{\ell,n}$ as one varies $\ell$ over the full range $\ell\geq 1$, with the above caveat of logarithmic interpretation of the power law when $\ell=1$. This is reminiscent of the behavior of models of statistical mechanics at the upper critical dimension (see [5] for a general introduction).
@@ 1.2. A brief outline of the proof
By definition, $Z_{\varnothing}^{[0]}=e^{-t}$, while $Z_{\alpha}^{[1]}(t)=\sum_{\delta=1}^{\infty}\delta^{\alpha-1}e^{\delta t}$ is a discrete analogue of $\int_{1}^{\infty}u^{\alpha-1}e^{-ut}{\rm d}u$ which is $t^{-\alpha}\Gamma(\alpha,t)$, in terms of the upper incomplete Gamma function. This explains the more convenient shift by 1 of the exponents for the $\delta$’s.
@@ 1.3. Relation to other work and possible directions for further inquiry
The original motivation for the present article was the log-concavity conjecture by the first author for the numbers $A(\ell,n,k)$ with respect to $k$ [1] (see also [22, Challenge 3] for $\ell=2$). This, of course, is equivalent to the log-concavity of the modified sequence $A(\ell,n,k)x^{k}$ which is proportional to the PMF of the random variables $\mathsf{K}_{\ell,n}$. Since the density of the standard Gaussian is the quintessential log-concave function, our CLT provides, in a weak sense, some confirmation for the conjecture. Moreover, it is the first such result which explores typical values for $k$ in the range $0\leq k\leq n$. The only proven results about the conjecture concern the “dilute polymer gas” regime where values of $k$ are very close to $n$, i.e., where nontrivial joint orbits are small and rare inside the environment $[n]$ (see [21, Cor. 4][2, Prop. 3.1], and [36]).
Obvious directions to explore beyond our CLT result are the study of moderate and large deviations for the $\mathsf{K}_{\ell,n}$, and upgrading the CLT to a local CLT. The latter would require asymptotics of the $A(\ell,n,k)$, rather than $\mathcal{H}_{\ell,n}(x)$, when both $n$ and $k$ are suitably large (see [32] for some steps in this direction). This would also be needed to prove the log-concavity conjecture in that range. Another interesting possible upgrade would be a functional CLT for a process associated to our $\mathbb{P}_{\ell,n,x}$ measures, in the spirit of [20] (see [30] for an in-depth study of such processes).
@@
Using the results of [2], it is easy to generalize the Bryan-Fulman formula, and our measures $\mathbb{P}_{\ell,n,x}$ so the weight per joint orbit $x$ can depend on the size or other finer features of the orbit. In the case of a single permutation, the only (conjugation) isomorphism invariant of an orbit, i.e., cycle is the size. For $\ell\geq 2$, the discrete $\ell$-tori have more “moduli” such as directional linear sizes, and a parameter (denoted $z$ in [2]) for the twist used when gluing the ends of a cylinder based on an $(\ell-1)$-dimensional similar discrete torus. This should give rise to more general measures than $\mathbb{P}_{\ell,n,x}$. Allowing only a dependence of $x$ on the size of an orbit is straightforward, since the $A(\ell,n,k)$ can be written in terms of the $A(\ell,n,1)$ (see [2, Thm. 1.1]), e.g., using Bell polynomials (see [30, Ch. 1]). This is also called a polymer gas representation in rigourous statistical mechanics [18].
@@
Finally, note that a random tuple of commuting permutations $(\sigma_{1},\ldots,\sigma_{\ell})$ is the same as a random group homomorphism $\varphi\in{\rm Hom}(\mathbb{Z}^{\ell},\mathfrak{S}_{n})$. There is a vast body of literature studying, more generally, random homomorphisms $\varphi\in{\rm Hom}(\Gamma,\mathfrak{S}_{n})$ where $\Gamma$ is a discrete finitely presented group (see, e.g., [25] and references therein). There, the emphasis is on groups that are far from commutative, unlike $\mathbb{Z}^{\ell}$, but the type of questions investigated is quite different from the focus of this article. Most of the results concern the uniform measure on ${\rm Hom}(\Gamma,\mathfrak{S}_{n})$. One can easily define analogues of our $\mathbb{P}_{\ell,n,x}$ in this setting by weighting a homomorphism $\varphi$ by $x^{\kappa_{n}(\varphi(\Gamma))}$. Whether these Ewens-like measures could be of use in the area of [25] remains to be seen. A good class of groups $\Gamma$ which may be worth looking at, from the CLT perspective of our article, are right-angled Artin groups which interpolate between $\mathbb{Z}^{\ell}$ and the free group on $\ell$ elements (see, e.g., [26]).
@@ 2. Preliminaries on the $Z$ functions
$\Delta(t)$ ::=:= $\int_{t}^{\infty}f(u)\ {\rm d}u-t\sum_{\delta=1}^{\infty}f(\delta t)$
@@
$Z_{\alpha_{1},\ldots,\alpha_{\ell}}^{[\ell]}(t)=\sum_{\delta_{1}=1}^{\infty}\delta_{1}^{\alpha_{1}-1}Z_{\alpha_{2},\ldots,\alpha_{\ell}}^{[\ell-1]}(t)=t^{-\alpha_{1}}\times t\sum_{\delta_{1}=1}^{\infty}f(\delta_{1}t)\ ,$
@@ 3. Preparation for the saddle point analysis
For $\0u\in\mathbb{C}\backslash(-\infty,0]$ we will use ${\rm Log}\ u:=\ln|u|+i{\rm Arg}\ u$ for the principal branch of the complex logarithm, namely, with $-\pi<{\rm Arg}\ u<\pi$. We of course have
@@
$\mathcal{L}_{\ell}(z)$ ::=:= $-\sum_{\delta_{1},\ldots,\delta_{\ell-1}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\ {\rm Log}(1-z^{\delta_{1}\cdots\delta_{\ell-1}})$ (17)
@@
Now that (5) is justified with $r=e^{-t}$ for some $t>0$ to be chosen shortly, we look at the modulus of the integrand $|z^{-n}\mathcal{G}_{\ell}(x,z)|=\exp(nt+x{\rm Re}\ \mathcal{L}_{\ell}(z))$. For $z=e^{-t+i\theta}$, with $-\pi<\theta<\pi$, we have
@@
$\Psi_{\ell,n}(s)$ ::=:= $\ln\mathbb{E}\left[\exp\left(s\left(\frac{\mathsf{K}_{\ell,n}-a_{n}}{b_{n}}\right)\right)\right]$ (22)
@@
where, for $y,u>0$ and $-\pi<\theta<\pi$, the $q$ functions are defined by
$q_{\ell,n}(y,u,\theta)$ ::=:= $in\theta-y\left(\mathcal{L}_{\ell}(e^{-u+i\theta})-\mathcal{L}_{\ell}(e^{-u})\right)$ (26)
$=$ $in\theta-y\sum_{\delta_{1},\ldots,\delta_{\ell}}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}e^{-\delta_{1}\cdots\delta_{\ell}u}\left(e^{i\delta_{1}\cdots\delta_{\ell}\theta}-1\right)\ .$
@@
${\rm Re}\ q_{\ell,n}(y,u,\theta)=2y\sum_{\delta_{1},\ldots,\delta_{\ell}}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}u}\ \sin^{2}\left(\frac{\delta_{1}\cdots\delta_{\ell}\theta}{2}\right)\ .$
@@ 4.2. Minor arc estimates
Note that for any fixed $c>1$, the function $v\mapsto\frac{1-v}{c^{2}-2cv+1}$ on the interval $(-\infty,\frac{c^{2}+1}{2c})$ which contains $[-1,1]$, has derivative $-\left(\frac{c-1}{c^{2}-2cv+1}\right)^{2}<0$. Therefore, ${\rm Re}\ \rho(u,\theta)$ is a decreasing function of $\cos\theta$ or rather $\cos|\theta|$, by parity. We now apply this to $y=x_{n}$, $u=t_{n}$, and we suppose $\theta\in\mathcal{R}_{\rm min}$, i.e., $\cos|\theta|\leq\cos t_{n}$, with $n$ large enough so that $t_{n}<\pi$. The last decreasing property then provides us with the lower bound
@@
$e^{2t_{n}}-2e^{t_{n}}\cos(t_{n})+1$ $=$ $(1+2t_{n}+2t_{n}^{2})-2\left(1+t_{n}+\frac{t_{n}^{2}}{2}\right)\left(1-\frac{t_{n}^{2}}{2}\right)+1+O(t_{n}^{3})$

### v2

@@ 1.1. The main result
For $n\geq 0$, the symmetric group $\mathfrak{S}_{n}$ comes with a natural left action $(\sigma,i)\mapsto\sigma(i)$ on the set $n]:=\{1,2,\ldots,n\}$. If $H$ is a subgroup of $\mathfrak{S}_{n}$, then the restriction of this action to $H$ gives us access to the enumerative quantity $\kappa_{n}(H):=|[n]/H|$, where we used $|\cdot|$ to denote the cardinality of finite sets, and where $[n]/H$ is the set of orbits for the action of $H$ on $[n]$. We will be particularly interested in the situation where $H=\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle$ is the subgroup generated by a random ordered $\ell$-tuple $(\sigma_{1},\ldots,\sigma_{\ell})$ of pairwise commuting permutations, for some fixed $\ell\geq 1$. Our main result is a central limit theorem for the number of joint orbits $\kappa_{n}(\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle)$ or simply $\kappa_{\ell,n}(\sigma_{1},\ldots,\sigma_{\ell})$. More precisely, let us introduce the set of commuting tuples of permutations
@@
We view $\mathbb{P}_{\ell,n,x}$ as a natural generalization of the Ewens measure with parameter $x$ to tuples of commuting permutations. Indeed, when $\ell=1$, $\kappa_{1,n}(\sigma)$ is just the number of cycles of the single permutation $\sigma$, and the well known expansion of the Pochhammer symbol in terms of Stirling numbers of the first kind gives the following specialization of the above PMF formula
@@
This is the PMF of the much studied Ewens distribution on the symmetric group, which was introduced in the applied context of mathematical biology [18], but is also an important notion from a pure mathematics point of view (see [35, 14, 40] for insightful reviews, and the book [6] for an in-depth introduction to the Ewens measure and many other related topics in probability theory and combinatorics). The uniform $x=1$ case was of course studied much earlier than the article by Ewens, but we also note that the particular weight $x=2$ per cycle (in a more difficult spatially dependent rather than mean field situation) features in the work of Tóth on quantum spin systems [42, §5].
@@
proved by Hansen as the time 1 projection of a functional CLT for a process related to the Ewens measure [25]. For $\ell=1$, the mean and variance are both asymptotically equivalent to $x\ln n$. The $x=1$ uniform sampling case is much older and due to Goncharov [21, 22] and is often given as an instructive example in graduate probability textbooks (see, e.g., [8, Example 27.3]). Using the Feller coupling idea [19, p. 815], one can express $\mathsf{K}_{1,n}$ as a sum of $n$ independent Bernoulli random variables of parameters $\frac{1}{n-j+1}$, $1\leq j\leq n$. Since these are not identically distributed, they provide a nice pedagogical example of application of the Lindeberg CLT. We also note the remarkable continuity of the formulas for the asymptotic mean and variance of the $\mathsf{K}_{\ell,n}$ as one varies $\ell$ over the full range $\ell\geq 1$, with the above caveat of logarithmic interpretation of the power law when $\ell=1$. This is reminiscent of the behavior of models of statistical mechanics at the upper critical dimension (see [7] for a general introduction).
@@ 1.2. A brief outline of the proof
By definition, $Z_{\varnothing}^{[0]}=e^{-t}$, while $Z_{\alpha}^{[1]}(t)=\sum_{\delta=1}^{\infty}\delta^{\alpha-1}e^{-\delta t}$ is a discrete analogue of $\int_{1}^{\infty}u^{\alpha-1}e^{-ut}{\rm d}u$ which is $t^{-\alpha}\Gamma(\alpha,t)$, in terms of the upper incomplete Gamma function. This explains the more convenient shift by 1 of the exponents for the $\delta$’s.
@@ 1.3. Relation to other work and possible directions for further inquiry
The original motivation for the present article was the log-concavity conjecture by the first author for the numbers $A(\ell,n,k)$ with respect to $k$ [1] (see also [27, Challenge 3] for $\ell=2$). This, of course, is equivalent to the log-concavity of the modified sequence $A(\ell,n,k)x^{k}$ which is proportional to the PMF of the random variables $\mathsf{K}_{\ell,n}$. Since the density of the standard Gaussian is the quintessential log-concave function, our CLT provides, in a weak sense, some confirmation for the conjecture. When $\ell\geq 3$, the only previously available results about the conjecture concern the “dilute polymer gas” regime where values of $k$ are very close to $n$, i.e., where nontrivial joint orbits are small and rare inside the environment $[n]$ (see [26, Cor. 4][3, Prop. 3.1], and [43]). Note that for $\ell=2$, the log-concavity was established for a wide range of values of $k$ when $n$ is large enough  [44].
Obvious directions to explore beyond our CLT result are the study of moderate and large deviations for the $\mathsf{K}_{\ell,n}$, and upgrading the CLT to a local CLT. The latter would require asymptotics of the $A(\ell,n,k)$, rather than $\mathcal{H}_{\ell,n}(x)$, when both $n$ and $k$ are suitably large (see [39, 2] for some results in this direction). Another interesting possible upgrade would be a functional CLT for a process associated to our $\mathbb{P}_{\ell,n,x}$ measures, in the spirit of [25] (see [36] for an in-depth study of such processes).
@@
Using the results of [3], it is easy to generalize the Bryan-Fulman formula, and our measures $\mathbb{P}_{\ell,n,x}$ so the weight per joint orbit $x$ can depend on the size or other finer features of the orbit. In the case of a single permutation, the only (conjugation) isomorphism invariant of an orbit, i.e., cycle is the size. For $\ell\geq 2$, the discrete $\ell$-tori have more “moduli” such as directional linear sizes, and a parameter (denoted $z$ in [3]) for the twist used when gluing the ends of a cylinder based on an $(\ell-1)$-dimensional similar discrete torus. This should give rise to more general measures than $\mathbb{P}_{\ell,n,x}$. Allowing only a dependence of $x$ on the size of an orbit is straightforward, since the $A(\ell,n,k)$ can be written in terms of the $A(\ell,n,1)$ (see [3, Thm. 1.1]), e.g., using Bell polynomials (see [36, Ch. 1]). This is also called a polymer gas representation in rigourous statistical mechanics [23].
@@
Note that a random tuple of commuting permutations $(\sigma_{1},\ldots,\sigma_{\ell})$ is the same as a random group homomorphism $\varphi\in{\rm Hom}(\mathbb{Z}^{\ell},\mathfrak{S}_{n})$. There is a vast body of literature studying, more generally, random homomorphisms $\varphi\in{\rm Hom}(\Gamma,\mathfrak{S}_{n})$ where $\Gamma$ is a discrete finitely presented group (see, e.g., [30] and references therein). There, the emphasis is on groups that are far from commutative, unlike $\mathbb{Z}^{\ell}$, but the type of questions investigated is quite different from the focus of this article. Most of the results concern the uniform measure on ${\rm Hom}(\Gamma,\mathfrak{S}_{n})$. One can easily define analogues of our $\mathbb{P}_{\ell,n,x}$ in this setting by weighting a homomorphism $\varphi$ by $x^{\kappa_{n}(\varphi(\Gamma))}$. Whether these Ewens-like measures could be of use in the area of [30] remains to be seen. A class of groups $\Gamma$ which may be worth looking at, from the CLT perspective of our article, are right-angled Artin groups which interpolate between $\mathbb{Z}^{\ell}$ and the free group on $\ell$ elements (see, e.g., [31]).
Finally, we would like to mention that our work is very closely related to the work of Ercolani and Ueltschi [17] as well as that of Maples, Nikeghbali and Zeindler [32]. Indeed, there has been much recent activity (see [16] and references therein) studying the so-called generalized Ewens measures, i.e., probability distributions on single permutations $\sigma\in\mathfrak{S}_{n}$, with probability mass function
$\mathbb{P}(\sigma)=\frac{1}{h_{n}n!}\prod_{j=1}^{n}\theta_{j}^{c_{j}(\sigma)}\ ,$
where $c_{j}(\sigma)$ is the number of cycles of length $j$ in the permutation $\sigma$, and the normalization factor is
$h_{n}=\frac{1}{n!}\sum_{\sigma\in\mathfrak{S}_{n}}\prod_{j=1}^{n}\theta_{j}^{c_{j}(\sigma)}\ .$
The measure thus depends on the choice of weights $(\theta_{j})_{j\geq 1}$. The basic Ewens measure corresponds to constant weights $\theta_{j}=x$. It is an easy exercise to show that our random variables $\mathsf{K}_{\ell,n}$ can be recovered as the number of cycles of the single random permutation $\sigma$ sampled according to such a generalized Ewens measure for the choice of weights
$\theta_{j}=\frac{x\ A(\ell,j,1)}{(j-1)!}\ .$
For $\ell\geq 2$, these are, on average, polynomially growing weights, i.e., belong to a less studied class of models in comparison to the case of logarithmically growing weights [6]. Nevertheless, there are some results on polynomially growing weights. For certain specific but different $\theta_{j}$ that on average grow like ours, mean asymptotics consistent with ours have been obtained in [17], while a CLT for these simpler polynomially growing weights was derived in [32]. Note that our weights grow like a power of $j$ only in an averaged or Cesaro sense. When $\ell=2$, we have $\theta_{j}=x\sigma(j)$ where $\sigma(j)=\sum_{q|j}q$ is the sum of divisors function of number theory. Simply showing sharp upper bounds on such a function is tantamount to the Riemann Hypothesis [37]. Note that it is possible that one may be able to prove our CLT using the general theorem in [32], but this would require showing the generating function for our $\theta_{j}$ is log-admissible à la Hayman. This typically is a nontrivial task (see, e.g., [13, Theorem 6.2]). Although exploring a different direction of generalization, one should also mention the work [4] on asymptotics for generating series of similar flavor to the one considered by Bryan and Fulman.
@@ 2. Preliminaries on the $Z$ functions
$\Delta(t)$ $:=$ $\int_{t}^{\infty}f(u)\ {\rm d}u-t\sum_{\delta=1}^{\infty}f(\delta t)$
@@
$Z_{\alpha_{1},\ldots,\alpha_{\ell}}^{[\ell]}(t)=\sum_{\delta_{1}=1}^{\infty}\delta_{1}^{\alpha_{1}-1}Z_{\alpha_{2},\ldots,\alpha_{\ell}}^{[\ell-1]}(\delta_{1}t)=t^{-\alpha_{1}}\times t\sum_{\delta_{1}=1}^{\infty}f(\delta_{1}t)\ ,$
@@ 3. Preparation for the saddle point analysis
For $u\in\mathbb{C}\backslash(-\infty,0]$ we will use ${\rm Log}\ u:=\ln|u|+i{\rm Arg}\ u$ for the principal branch of the complex logarithm, namely, with $\pi<{\rm Arg}\ u<\pi$. We of course have
@@
$\mathcal{L}_{\ell}(z)$ $:=$ $-\sum_{\delta_{1},\ldots,\delta_{\ell-1}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\ {\rm Log}(1-z^{\delta_{1}\cdots\delta_{\ell-1}})$ (17)
@@
Now that (5) is justified with $r=e^{-t}$ for some $t>0$ to be chosen shortly, we look at the modulus of the integrand $|z^{-n}\mathcal{G}_{\ell}(x,z)|=\exp(nt+x{\rm Re}\ \mathcal{L}_{\ell}(z))$. For $z=e^{-t+i\theta}$, with $\pi<\theta<\pi$, we have
@@
$\Psi_{\ell,n}(s)$ $:=$ $\ln\mathbb{E}\left[\exp\left(s\left(\frac{\mathsf{K}_{\ell,n}-a_{n}}{b_{n}}\right)\right)\right]$ (22)
@@
where, for $y,u>0$ and $\pi<\theta<\pi$, the $q$ functions are defined by
$q_{\ell,n}(y,u,\theta)$ $:=$ $in\theta-y\left(\mathcal{L}_{\ell}(e^{-u+i\theta})-\mathcal{L}_{\ell}(e^{-u})\right)$ (26)
$=$ $in\theta-y\sum_{\delta_{1},\ldots,\delta_{\ell}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}e^{-\delta_{1}\cdots\delta_{\ell}u}\left(e^{i\delta_{1}\cdots\delta_{\ell}\theta}-1\right)\ .$
@@
${\rm Re}\ q_{\ell,n}(y,u,\theta)=2y\sum_{\delta_{1},\ldots,\delta_{\ell}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}u}\ \sin^{2}\left(\frac{\delta_{1}\cdots\delta_{\ell}\theta}{2}\right)\ .$
@@ 4.2. Minor arc estimates
Note that for any fixed $c>1$, the function $v\mapsto\frac{1-v}{c^{2}-2cv+1}$ on the interval $(-\infty,\frac{c^{2}+1}{2c})$ which contains [1,1][-1,1], has derivative $\left(\frac{c-1}{c^{2}-2cv+1}\right)^{2}<0$. Therefore, ${\rm Re}\ \rho(u,\theta)$ is a decreasing function of $\cos\theta$ or rather $\cos|\theta|$, by parity. We now apply this to $y=x_{n}$, $u=t_{n}$, and we suppose $\theta\in\mathcal{R}_{\rm min}$, i.e., $\cos|\theta|\leq\cos t_{n}$, with $n$ large enough so that $t_{n}<\pi$. The last decreasing property then provides us with the lower bound
@@
$e^{2t_{n}}-2e^{t_{n}}\cos(t_{n})+1$ $=$ $(1+2t_{n}+2t_{n}^{2})-2\left(1+t_{n}+\frac{t_{n}^{2}}{2}\right)\left(1-\frac{t_{n}^{2}}{2}\right)+1+O(t_{n}^{3})\$